Identify the type of the conic 4(x-2y+1)^2 +9(2x+y+2)^2 = 25
1
Expert's answer
2020-11-24T15:26:34-0500

Given "4(x-2y+1)^2 + 9(2x+y+2)^2" "= 25"

Divide through by 25, this gives us

"\frac{(x-2y+1)^2}{\frac{25}{4}} + \frac{(2x+y+2)^2}{\frac{25}{9}}"

"=\frac{(x-(2y-1))^2}{(\frac{5}{2})^2} + \frac{((y-(-2x-2))^2}{(\frac{5}{3})^2}=1"

which satisfies the equation of ellipse

"\frac{(x1)^2}{a^2} + \frac{(y1)^2}{b^2}=1" ,

where "x1, y1" are new variables.

Therefore the conic is an ellipse.